OpenStudy (anonymous):

c=5(f-32)/9 solve for f

OpenStudy (akashdeepdeb):

\[c = \frac{5}{9} (f - 32)\] Multiply both sides by \(\frac{9}{5}\). \[\frac{9}{5} c = (f - 32) \] Add \(32\) on both sides. \[\frac{9}{5} c + 32 = f\] or \[f = \frac{9}{5} c + 32\] Understood this? :)